RC Meeting and Run Plan, Fri, January 24th 2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

General updates

  • Silviu posted a simulation of glass cell window response to raster breathing rate [1]
  • JP posted the procedure for moving from transverse to longitudinal polarization [2]

Thursday swing – Friday owl

  • 4:30PM: Cryo problems (which caused Thursday's magnet problems) resolved
  • 5:00PM – 8PM: Beam restoration
  • 8PM – midnight: Took about 4 hours of good transverse data
  • Midnight: Arc dipole trips again.
    • Resolved at 8AM Friday
    • PD says this was a loose wire, which has been taken care of [3]

Thursday day

  • Several SHMS dipole trips during Day Shift
    • Steve Lassiter said PSU hasn't failed "enough" to really diagnose what's wrong
    • So for now all we can do is follow procedure [4] [5] and hope for the best
  • MCC had some trouble getting width even below 500 um during Friday Day shift, but we ended up at 469um
  • MCC seems to have some trouble delivering 30uA to us at the same time as Hall A's 150 uA
  • Injector work in the morning
  • 1PM Start beam checkout
  • 3PM Taking transverse polarization data

Recent polarization measurements

 Trans  51%  2PM Thu [6]
 Trans  61%  7PM Thu [7]
 Trans  63% 12AM Fri [8]
 Trans  66%  6AM Fri [9]
 Long   63% 10AM Fri [10]
 Long   57%  2PM Fri [11]

Beam checkout

  • I wrote a script for projecting beam width to target [12]
    • The distances Simona used in her post [13] were from CED. Apparently CED is not updated with the most recent survey results. We will ask Jay to update this.
    • Thanks to Dave Gaskell for updated distances [14]
  • I'm updating the short checkout procedure wiki
    • To be run any time something(?) changes in the beam line
    1. Harp scan with tune beam and no target
      • My script is helpful here
    2. YAG screenshot with and without raster, tune(?) beam, no target
      • Verify raster off spot looks round (not like a comet, no weird tails)
      • Verify raster on beam looks like a circle
      • This will interrupt the other halls' beam
    3. Check beam positions with 30 uA CW beam, no target. Make sure orbit lock and current ramp lock are on.
    4. Mask FSD, move target, start taking data

Current Run Plan

  • Always take 120Hz of EDTM


!!                                                     !!
!!  All NMR measurement HCLOG entries should include:  !!
!!     1) A title that clearly lists polarization      !!
!!     2) Screenshots of both NMR_FieldSweep and       !!
!!        the polarization fit.                        !!
!!                                                     !!
!!  For an example see the following entry:            !!
!!     https://logbooks.jlab.org/entry/3768696       !!
!!                                                     !!
  • Polarized 3He target
    • We're taking transverse data now
    • And over the weekend?

Moller & spin dance

  • Reminder: we only have one measurement so far from Saturday 1/18
  • Tentatively Monday morning, pending Hall A
  • Hall B strongly encourages us to push it back to the 30th (because it won't matter to them at that point and they currently have priority)


  • Open enclosure early next week, maybe Tuesday 1/28
    • Check longitudinal laser setup
    • Install the rotation stages?
    • Install EPR fiber bundle?
    • Check to see if there's any scheduled downtime that we can use for this
  • Arun tentatively getting replacement cell prepared early next week; install it in two(?) weeks?

Special Requests

  • Keep target expert schedule filled out on whiteboard

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
  • Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
  • The operators have been somewhat careful with our beam orbit locks. Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
  • Ramp rate lock instructions are here and added to checklist[15]
  • Temporary Laser instructions can be found here [16]

Planned Accelerator Configuration Changes